Solve the differential equation ( 2y - y + 1)dx + ( 2y - x - 1) dy.

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Rakesh
 
I am answering it as original question because you have done some mistake
 
 
So we can write it as 2xdx+dx+2ydy-dy= xdy+ydx
But d(xy)= xdy+ydx
So 2xdx+dx+2ydy-dy= dxy
Integrate both sides
x^2+x+y^2-y= xy+C